Suppose there are only two individuals in the market for some product. Individual A’s inverse demand equation is P = 8 - 0.5 QD and B’s inverse demand equation is P = 10 - QD

a) Derive the equation showing A’s quantity demanded as a function of price. Ie: A’s demand equation

b) Derive B’s demand equation

c) Derive the aggregate (market) inverse demand equation. Ie: the equation showing price as a function of aggregate demand

d) Derive the aggregate (market) demand equation
1
Expert's answer
2020-01-27T10:11:40-0500

a) A's demand equation

"P = 8 - 0.5QD."

"P-8=-0.5QD"

"\frac{P-8}{-0.5}=QD"

"-2P+16=QD"

Therefore, A's demand equation is "QD=16-2P"

b) B's demand equation

"P = 10 - QD"

"P-10=-QD"

"-P+10=QD"

Thus, B's demand equation is "QD=10-P"

c) since demand cannot be negative, A enters the market when price is less than 8 and B when price is less than 10. Therefore, when price is between 8 and 10 the aggregate inverse demand is given by "P=10-QD" (B's inverse demand equation). But when price is less than 8, the aggregate inverse demand equation is given by;

"\frac{P-8}{-0.5}=QD+P-10"

"P-8=-0.5QD-0.5P+5"

"1.5P=13-0.5QD"

"P=\frac{26-QD}{3}"

d) the aggregate demand when price is between 8 and 10 is given by "QD=10-P" (B's demand equation). But when price is less than 8, the aggregate demand is given by "QD=10-P+16-2P"

"QD=26-3p"